# Subvarieties of pointed Abelian $\ell$-groups

### v1

@@ 3 The lexicographic decompositions
This section focuses on understanding the structure of the lexicographic product of $\ell$-groups. The lexicographic product is a key operation on $\ell$-groups. It allows for the creation of otherwise unintuitive $\ell$-groups that play an essential role in various classifications (see e.g. the famous Hahn Theorem from [18]). Since we will need to use this tool frequently in the following chapters, here we establish several basic properties of this construction. The main result of this section is Theorem 3.10, which is a stronger version of [28, Lemma 4.6], which tells us that from the viewpoint of universal classes of totally ordered Abelian $\ell$-groups we are not interested in $\ell$-groups without a strong unit. We will proceed to the definition of the lexicographic product.
@@
Note that for totally ordered Abelian $\ell$-groups ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}}$ and Abelian $\ell$-group $\boldsymbol{\mathit{C}}$} it holds that ${\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ is totally ordered and moreover $({\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{C}}}\cong{\boldsymbol{\mathit{A}}}\overrightarrow{\times}({\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{C}}})$. Therefore, we will commonly omit parentheses and we will just write ${\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{C}}}$.
@@
Since $\boldsymbol{\mathit{A}}$} is finitely generated, also ${\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ is finitely generated. Since ${\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ is a finitely generated torsion-free Abelian group, it is group-isomorphic (but not necessarily $\ell$-group isomorphic) to a free Abelian group (see [27, Theorem 10.19]). Thus the group exact sequence $0\rightarrow{\boldsymbol{\mathit{B}}}\rightarrow^{\iota}{\boldsymbol{\mathit{A}}}\rightarrow^{\pi}{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow 0$ splits and there is a group homomorphism $p:{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}$ such that $\pi\circ p=id_{{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}}$. It is well-known (see [27, Lemma 10.3]), that the mapping $\varphi:({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\times{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}$ defined as $\langle{x,y}\rangle\mapsto p(x)+\iota(y)$ is a group isomorphism. We want to show that $\varphi$ is an $\ell$-group isomorphism between $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ and $\boldsymbol{\mathit{A}}$} as well. We will show that $\varphi$ is order preserving. Let us pick $\langle{a,b}\rangle,\langle{c,d}\rangle\in(A/B)\overrightarrow{\times}B$ such that $\langle{a,b}\rangle\leq\langle{c,d}\rangle$. We distinguish two cases:
@@
    If $a<c$ we get $\pi(p(a)+\iota(b))<\pi(p(c)+\iota(d))$ since
@@
    If $a=c$ then $b\leq d$ so clearly $p(a)+\iota(b)\leq p(c)+\iota(d)$.
This proves that $\varphi$ is an $\ell$-group homomorphism. It remains to prove that $\varphi^{-1}$ preserves ordering as well. This follows easily, from the fact that $\boldsymbol{\mathit{A}}$} is a totally ordered algebra and $\varphi$ is an order preserving isomorphism. Thus $\varphi$ is $\ell$-group isomorphism between $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ and $\boldsymbol{\mathit{A}}$}. ∎
One can easily prove a pointed version of this lemma, by just verifying, that the isomorphism in the proof of the Lemma 3.2 preserves the point structure.
@@
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a finitely generated totally ordered pointed Abelian $\ell$-group and ${\boldsymbol{\mathit{B}}}_{b}$ be a convex pointed $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{b}$. Then ${\boldsymbol{\mathit{A}}}_{b}\cong({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})_{0}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}$.
@@
Let us denote a strong unit of $\boldsymbol{\mathit{A}}$} by $b$. Consider an ultrapower of ${\boldsymbol{\mathit{A}}}_{a}$ defined as ${\boldsymbol{\mathit{C}}}=\prod_{i\in\omega}{\boldsymbol{\mathit{A}}}_{a}/\mathfrak{U}$, where $\mathfrak{U}$ is a non-principal ultrafilter on $\omega$ and let $\iota:d\mapsto\langle{d,\dots,d}\rangle$ be the canonical embedding of $\boldsymbol{\mathit{A}}$} into $\boldsymbol{\mathit{C}}$}. Consider the element $c\in C$ defined as $c=\langle{b,2b,3b,\dots}\rangle$. Clearly, $c>\iota(d)$ for any $d\in{\boldsymbol{\mathit{A}}}$. Let us denote by $\boldsymbol{\mathit{B}}$} the $\ell$-subgroup of $\boldsymbol{\mathit{C}}$} generated by $\iota[A]\cup\{c\}$. For any $d\in A$ we have $c<\iota(d)<c$ thus we can uniquely express any element of $B$ as $n\cdot c+\iota(d)$ for some $n\in\mathbb{Z}$ and $d\in A$. For any $n_{1},n_{2}\in\mathbb{Z}$ and $d_{1},d_{2}\in A$ we have $n_{1}\cdot c+\iota(d_{1})\geq n_{2}\cdot c+\iota(d_{2})$ iff $n_{1}>n_{2}$ or $n_{1}=n_{2}$ and $d_{1}\geq d_{2}$. This proves ${\boldsymbol{\mathit{B}}}_{\iota(a)}\cong{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}$. Since ${\boldsymbol{\mathit{B}}}_{\iota(a)}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{a})$, we conclude ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{a})$. ∎
@@
Let $\mathbb{K}$}} be a nontrivial subvariety of $p\mathbb{AL}$. Take any ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}$. Take the subalgebra of ${\boldsymbol{\mathit{A}}}_{a}$ generated by $a$. Such an algebra has to be isomorphic to ${\boldsymbol{\mathit{Z}}}_{1}$, ${\boldsymbol{\mathit{Z}}}_{0}$ or ${\boldsymbol{\mathit{Z}}}_{-1}$. If we get ${\boldsymbol{\mathit{Z}}}_{1}\in{{\mathbb{K}}}$ or ${\boldsymbol{\mathit{Z}}}_{-1}\in{{\mathbb{K}}}$, by Lemma 3.5 we get ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{1}\in{{\mathbb{K}}}$ or ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{-1}\in{{\mathbb{K}}}$. Since $\mathbb{K}$}} is closed under $\mathbf{H}$, we get ${\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}$ in all cases. Therefore, ${\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}$ and by Lemma 2.11 we conclude that $p\mathbb{AL}^{0}\subseteq{{\mathbb{K}}}$. ∎
@@
Let ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}},{\boldsymbol{\mathit{C}}},{\boldsymbol{\mathit{D}}}$ be $\ell$-group chains and assume $\boldsymbol{\mathit{A}}$} partially embeds into $\boldsymbol{\mathit{B}}$}. Then ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$ partially embeds into ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$.
@@
Therefore, $\psi_{G}$ preserves addition, subtraction and zero constant. It remains to check $\psi_{G}$ preserves operations $\lor,\land$. Since ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}},{\boldsymbol{\mathit{C}}},{\boldsymbol{\mathit{D}}}$ are chains it is enough to show that $\psi_{G}$ preserves the lattice ordering. Let us assume that $\langle{c_{1},a_{1},d_{1}}\rangle,\langle{c_{2},a_{2},d_{2}}\rangle\in G$ and $\langle{c_{1},a_{1},d_{1}}\rangle<\langle{c_{2},a_{2},d_{2}}\rangle$. Let us distinguish three cases:
@@
We can easily observe that the lemma also holds for pointed $\ell$-groups.
@@
Let ${\boldsymbol{\mathit{A}}}_{a},{\boldsymbol{\mathit{B}}}_{b},{\boldsymbol{\mathit{C}}}_{c},{\boldsymbol{\mathit{D}}}_{d}$ be pointed $\ell$-group chains and assume ${\boldsymbol{\mathit{A}}}_{a}$ partially embeds into ${\boldsymbol{\mathit{B}}}_{b}$. Then ${\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$ partially embeds into ${\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$.
@@
Finally, we conclude this section with the following theorem which is a strengthening of [28, Lemma 4.6].
@@
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a finitely generated totally ordered pointed Abelian $\ell$-group and ${\boldsymbol{\mathit{B}}}_{b}$ be its convex strongly pointed $\ell$-subgroup, with $b\neq 0$. [^4] Let us note, that that such ${\boldsymbol{\mathit{B}}}_{b}$ is the smallest pointed convex $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{b}$. As such, ${\boldsymbol{\mathit{B}}}_{b}$ is unique as it is uniquelly determined by ${\boldsymbol{\mathit{A}}}_{b}$. Then $\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{b})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{B}}}_{b})$.
@@
Theorem 3.10 tells us that the universal theory of any totally ordered pointed $\ell$-group is equal to the universal theory generated by its convex $\ell$-subgroup, which is strongly pointed or the universal theory generated by ${\boldsymbol{\mathit{Z}}}_{0}$. In other words, when classifying universal classes of totally ordered pointed $\ell$-groups, we can restrict our focus only on those which are strongly pointed or 0-pointed.
4 Characterization of varieties generated by a single finitely generated totally ordered $\ell$-group
In this section we describe all join-irreducible subvarieties of pointed Abelian $\ell$-groups. Although, one could obtain the result of this chapter using Theorem 3.10 from Section 3, Mundici functor and Komori classification of MV-algebras, we have chosen a different, possibly harder approach, by providing a more self-contained theory and not using the theory of MV-algebras. However, a reader familiar with Komori classification of MV-algebras will find some of the proofs here possibly familiar, since they often use similar techniques (for comparison see [7]).
@@ 4 Characterization of varieties generated by a single finitely generated totally ordered $\ell$-group
We check that $\varphi$ preserves the lattice operations of lexicographic product.
@@
This shows $\varphi$ preserves $\lor$. Similarly we can show that $\varphi$ preserves $\land$. Therefore, $\varphi$ is an injective homomorphism.
In the case, when $\psi$ is an isomorphism, there exists the inverse isomorphism $\psi^{-1}:{\boldsymbol{\mathit{B}}}_{b}\rightarrow{\boldsymbol{\mathit{A}}}_{a}$. By the previous part of the proof $\varphi^{-1}:{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}\rightarrow{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$ defined as $\langle{x,y,z}\rangle\mapsto\langle{x,\psi^{-1}(y),z}\rangle$ is an injective homomorphism. Clearly, $\varphi^{-1}$ is also inverse to $\varphi$, which shows $\varphi$ has to be an isomorphism. ∎
@@
Let $a_{1},\dots,a_{m}\in\mathbb{R}$ and let $f:({\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}})^{m}\rightarrow{\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}$ be a pointed $\ell$-group term function. Let $\pi_{2}:{\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}\rightarrow{\boldsymbol{\mathit{R}}}$ be a projection to the second coordinate. Then the $m$-ary function
@@
is a continuous function from ${\boldsymbol{\mathit{R}}}^{m}$ to $\boldsymbol{\mathit{R}}$}. In particular, every pointed $\ell$-group term function on $\boldsymbol{\mathit{R}}$} is continuous.
@@
We show that $g$ is a composition of continuous functions. Clearly, constants are continuous functions. Moreover, addition and subtraction are defined component-wise thus they are continuous functions.
It remains to check maximum and minimum. For each constants $b_{1},b_{2}\in\mathbb{R}$ we define
@@
Thus for all $b_{1},b_{2}\in\mathbb{R}$ we obtain that $\lor_{b_{1},b_{2}}$ and $\land_{b_{1},b_{2}}$ are continuous functions. Since $g$ is a composition of addition, subtraction, constants and functions $\lor_{b_{i},b_{j}}$ and $\land_{b_{i},b_{j}}$ for some $b_{i},b_{i}\in\mathbb{R}$, we obtain that $g$ is continuous as well. ∎
The following lemma is the generalization of [7, Propositions 8.1.1 and 8.1.2]

### v2

@@ 3 The lexicographic decompositions
This section focuses on understanding the structure of the lexicographic product of $\ell$-groups. The lexicographic product is a key operation on $\ell$-groups. It allows for the creation of otherwise unintuitive $\ell$-groups that play an essential role in various classifications (see e.g. the famous Hahn Theorem from [20]). Since we will need to use this tool frequently in the following sections, here we establish several basic properties of this construction. The main result of this section is Theorem 3.10, which is a stronger version of [32, Lemma 4.6], which demonstrates that when classifying universal classes of totally ordered pointed Abelian $\ell$-groups, we can restrict our focus to those that are strongly pointed or 0-pointed. We will proceed to the definition of the lexicographic product.
@@
Note that for totally ordered Abelian $\ell$-groups ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}}$ and an Abelian $\ell$-group $\boldsymbol{\mathit{C}}$} it holds that ${\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ is totally ordered and moreover $({\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{C}}}\cong{\boldsymbol{\mathit{A}}}\overrightarrow{\times}({\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{C}}})$. Therefore, we will commonly omit parentheses and we will just write ${\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{C}}}$.
@@
Since $\boldsymbol{\mathit{A}}$} is finitely generated, ${\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ is also finitely generated. Since ${\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ is a finitely generated torsion-free Abelian group, it is group-isomorphic (but not necessarily $\ell$-group isomorphic) to a free Abelian group (see [31, Theorem 10.19]). Thus the group exact sequence $0\rightarrow{\boldsymbol{\mathit{B}}}\rightarrow^{\iota}{\boldsymbol{\mathit{A}}}\rightarrow^{\pi}{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow 0$ splits and there is a group homomorphism $p:{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}$ such that $\pi\circ p=id_{{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}}$. It is well-known (see [31, Lemma 10.3]) that the mapping $\varphi:({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\times{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}$ defined as $\langle{x,y}\rangle\mapsto p(x)+\iota(y)$ is a group isomorphism. We want to show that $\varphi$ is an $\ell$-group isomorphism between $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ and $\boldsymbol{\mathit{A}}$} as well. We will show that $\varphi$ is order-preserving. Let us pick $\langle{a,b}\rangle,\langle{c,d}\rangle\in(A/B)\times B$ such that $\langle{a,b}\rangle\leq\langle{c,d}\rangle$. We distinguish two cases:
@@
    If $a<c$, we obtain $\pi(p(a)+\iota(b))<\pi(p(c)+\iota(d))$ since
@@
    If $a=c$, then $b\leq d$, so clearly $p(a)+\iota(b)\leq p(c)+\iota(d)$.
This proves that $\varphi$ is an $\ell$-group homomorphism. It remains to prove that $\varphi^{-1}$ preserves ordering as well. This follows easily from the fact that $\boldsymbol{\mathit{A}}$} is a totally ordered algebra and $\varphi$ is an order preserving isomorphism. Thus $\varphi$ is an $\ell$-group isomorphism between $({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})\overrightarrow{\times}{\boldsymbol{\mathit{B}}}$ and $\boldsymbol{\mathit{A}}$}. ∎
Generally, it is well known that one can define a lexicographic ordering on a totally ordered Abelian group $\boldsymbol{\mathit{A}}$}, provided one has a short exact sequence of Abelian groups $0\rightarrow{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}\rightarrow{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow 0$ and ${\boldsymbol{\mathit{B}}},{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ are totally ordered [10, Problem 1.8]. However, one still needs to assume that ${\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}$ is finitely generated to prove that the sequence $0\rightarrow{\boldsymbol{\mathit{B}}}\rightarrow{\boldsymbol{\mathit{A}}}\rightarrow{\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}}\rightarrow 0$ splits. Let us also remark that one can easily prove a pointed version of this lemma by just verifying that the isomorphism in the proof of Lemma 3.2 preserves the point structure.
@@
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a finitely generated totally ordered pointed Abelian $\ell$-group and ${\boldsymbol{\mathit{B}}}_{b}$ be a pointed convex $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{b}$. Then ${\boldsymbol{\mathit{A}}}_{b}\cong({\boldsymbol{\mathit{A}}}/{\boldsymbol{\mathit{B}}})_{0}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}$.
@@
Let us note that for any pointed Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{b}$ with $b\neq 0$, there is a unique strongly pointed convex $\ell$-subgroup ${\boldsymbol{\mathit{B}}}_{b}$, where $\boldsymbol{\mathit{B}}$} is the $\ell$-subgroup of $\boldsymbol{\mathit{A}}$} generated (as a convex $\ell$-subgroup) by $b$, expanded by $b$ as its designated element.
@@
Let us denote a strong unit of $\boldsymbol{\mathit{A}}$} by $b$. Consider an ultrapower of ${\boldsymbol{\mathit{A}}}_{a}$ defined as ${\boldsymbol{\mathit{C}}}=\prod_{i\in\omega}{\boldsymbol{\mathit{A}}}_{a}/\mathfrak{U}$, where $\mathfrak{U}$ is a non-principal ultrafilter on $\omega$ and let $\iota:d\mapsto\langle{d,\dots,d}\rangle$ be the canonical embedding of $\boldsymbol{\mathit{A}}$} into $\boldsymbol{\mathit{C}}$}. Consider the element $c\in C$ defined as $c=\langle{|b|,2|b|,3|b|,\dots}\rangle$. Since the set $\{n\cdot|b|\geq|d|\}$ is cofinite for any $d\in{\boldsymbol{\mathit{A}}}$, we have $\{n\mid n\cdot|b|\geq|d|\}\in\mathfrak{U}$ and thus $c>|\iota(d)|$ for any $d\in{\boldsymbol{\mathit{A}}}$. Let us denote by $\boldsymbol{\mathit{B}}$} the $\ell$-subgroup of $\boldsymbol{\mathit{C}}$} generated by $\iota[A]\cup\{c\}$. For any $d\in A$ we have $c<\iota(d)<c$ thus we can uniquely express any element of $B$ as $n\cdot c+\iota(d)$ for some $n\in\mathbb{Z}$ and $d\in A$. For any $n_{1},n_{2}\in\mathbb{Z}$ and $d_{1},d_{2}\in A$ we have $n_{1}\cdot c+\iota(d_{1})\geq n_{2}\cdot c+\iota(d_{2})$ iff $n_{1}>n_{2}$ or $n_{1}=n_{2}$ and $d_{1}\geq d_{2}$. This proves ${\boldsymbol{\mathit{B}}}_{\iota(a)}\cong{\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}$. Since ${\boldsymbol{\mathit{B}}}_{\iota(a)}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{a})$, we conclude ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\in\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{a})$. ∎
@@
Let $\mathbb{K}$}} be a nontrivial subvariety of $p\mathbb{AL}$. Take any non-trivial totally ordered ${\boldsymbol{\mathit{A}}}_{a}\in{{\mathbb{K}}}$. Then ${\boldsymbol{\mathit{A}}}_{a}$ contains a pointed $\ell$-subgroup isomorphic to ${\boldsymbol{\mathit{Z}}}_{1}$, ${\boldsymbol{\mathit{Z}}}_{0}$ or ${\boldsymbol{\mathit{Z}}}_{-1}$. If we get ${\boldsymbol{\mathit{Z}}}_{1}\in{{\mathbb{K}}}$ or ${\boldsymbol{\mathit{Z}}}_{-1}\in{{\mathbb{K}}}$, by Lemma 3.5 we get ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{1}\in{{\mathbb{K}}}$ or ${\boldsymbol{\mathit{Z}}}_{0}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{-1}\in{{\mathbb{K}}}$. Since $\mathbb{K}$}} is closed under $\mathbf{H}$, we get ${\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}$ in all cases. Therefore, ${\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}$ and by Lemma 2.13 we conclude that $p\mathbb{AL}^{0}\subseteq{{\mathbb{K}}}$. ∎
@@
Let ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}},{\boldsymbol{\mathit{C}}},{\boldsymbol{\mathit{D}}}$ be totally ordered Abelian $\ell$-groups and assume $\boldsymbol{\mathit{A}}$} partially embeds into $\boldsymbol{\mathit{B}}$}. Then ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$ partially embeds into ${\boldsymbol{\mathit{C}}}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}$.
@@
Therefore, $\psi_{G}$ preserves addition, subtraction and zero constant. It remains to check $\psi_{G}$ preserves operations $\lor,\land$. Since ${\boldsymbol{\mathit{A}}},{\boldsymbol{\mathit{B}}},{\boldsymbol{\mathit{C}}},{\boldsymbol{\mathit{D}}}$ are totally ordered, it is enough to show that $\psi_{G}$ preserves the lattice ordering. Let us assume that $\langle{c_{1},a_{1},d_{1}}\rangle,\langle{c_{2},a_{2},d_{2}}\rangle\in G$ and $\langle{c_{1},a_{1},d_{1}}\rangle<\langle{c_{2},a_{2},d_{2}}\rangle$. Let us distinguish three cases:
@@
We can easily observe that the lemma also holds for pointed Abelian $\ell$-groups.
@@
Let ${\boldsymbol{\mathit{A}}}_{a},{\boldsymbol{\mathit{B}}}_{b},{\boldsymbol{\mathit{C}}}_{c},{\boldsymbol{\mathit{D}}}_{d}$ be totally ordered pointed Abelian $\ell$-groups and assume ${\boldsymbol{\mathit{A}}}_{a}$ partially embeds into ${\boldsymbol{\mathit{B}}}_{b}$. Then ${\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$ partially embeds into ${\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$.
@@
Finally, we conclude this section with the following theorem which is a strengthening of [32, Lemma 4.6].
@@
Let ${\boldsymbol{\mathit{A}}}_{b}$ be a finitely generated totally ordered pointed Abelian $\ell$-group with $b\neq 0$ and ${\boldsymbol{\mathit{B}}}_{b}$ be its strongly pointed convex $\ell$-subgroup. Then $\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{A}}}_{b})=\mathbf{ISP}_{\mathrm{U}}({\boldsymbol{\mathit{B}}}_{b})$.
@@
Theorem 3.10 tells us that the universal theory of any totally ordered pointed Abelian $\ell$-group is equal to the universal theory generated by its strongly pointed convex $\ell$-subgroup or the universal theory generated by ${\boldsymbol{\mathit{Z}}}_{0}$. In other words, when classifying universal classes of totally ordered pointed Abelian $\ell$-groups, we can restrict our focus to those which are strongly pointed or 0-pointed.
From Theorem 3.10 and Corollary 3.6 we obtain the following two additional corollaries.
Corollary 3.11.
Every non-trivial subvariety of $p\mathbb{AL}$ that is not equal to $p\mathbb{AL}^{0}$ is generated by its finitely generated totally ordered strongly pointed members.
Corollary 3.12.
Every non-trivial subquasivariety of $p\mathbb{AL}$ generated by its totally ordered members is also generated by those of its finitely generated totally ordered members that are either strongly pointed or 0-pointed.
4 Characterization of varieties generated by a single finitely generated totally ordered pointed Abelian $\ell$-group
In this section we describe all join-irreducible subvarieties of pointed Abelian $\ell$-groups. Although one could obtain the result of this section using Theorem 3.10 from Section 3, the Mundici functor and the Komori classification, we have chosen a different, more self-contained approach that does not rely on the theory of MV-algebras. However, a reader familiar with Komori classification of MV-algebras will find some of the proofs here possibly familiar, since they often use similar techniques (for comparison see [8]).
@@
We check that $\varphi$ preserves the lattice operations of the lexicographic product.
@@
This shows $\varphi$ preserves $\lor$. Similarly, we can show that $\varphi$ preserves $\land$. Therefore, $\varphi$ is an injective homomorphism.
In the case when $\psi$ is an isomorphism, there exists the inverse isomorphism $\psi^{-1}:{\boldsymbol{\mathit{B}}}_{b}\rightarrow{\boldsymbol{\mathit{A}}}_{a}$. By the previous part of the proof $\varphi^{-1}:{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{B}}}_{b}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}\rightarrow{\boldsymbol{\mathit{C}}}_{c}\overrightarrow{\times}{\boldsymbol{\mathit{A}}}_{a}\overrightarrow{\times}{\boldsymbol{\mathit{D}}}_{d}$ defined as $\langle{x,y,z}\rangle\mapsto\langle{x,\psi^{-1}(y),z}\rangle$ is an injective homomorphism. Clearly, $\varphi^{-1}$ is also inverse to $\varphi$, which shows $\varphi$ must be an isomorphism. ∎
@@
Let $a_{1},\dots,a_{m}\in\mathbb{R}$ and let $f:({\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}})^{m}\rightarrow{\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}$ be a pointed Abelian $\ell$-group term function. Let $\pi_{2}:{\boldsymbol{\mathit{R}}}\overrightarrow{\times}{\boldsymbol{\mathit{R}}}\rightarrow{\boldsymbol{\mathit{R}}}$ be a projection to the second coordinate. Then the $m$-ary function
@@
is a continuous function from ${\boldsymbol{\mathit{R}}}^{m}$ to $\boldsymbol{\mathit{R}}$}. In particular, every pointed Abelian $\ell$-group term function on $\boldsymbol{\mathit{R}}$} is continuous.
@@
We show that $g$ is a composition of continuous functions. We proceed by induction on the complexity of the term $f$. Clearly, constants and projections are continuous functions. Moreover, addition and subtraction are defined component-wise, thus they are continuous functions.
It remains to check maximum and minimum. For any constants $b_{1},b_{2}\in\mathbb{R}$ we define
@@
Thus for all $b_{1},b_{2}\in\mathbb{R}$ we obtain that $\lor_{b_{1},b_{2}}$ and $\land_{b_{1},b_{2}}$ are continuous functions. Since $g$ is a composition of addition, subtraction, constants and functions $\lor_{b_{i},b_{j}}$ and $\land_{b_{i},b_{j}}$ for some $b_{i},b_{j}\in\mathbb{R}$, we obtain that $g$ is continuous as well. ∎
The following lemma is the generalization of [8, Propositions 8.1.1 and 8.1.2]